Find the delta G degree for the following reaction
Answer #1

dG = dH - T*dS

dG = dGproducts - dGreactants

dG = (2*PCl5 + 5*H2) - (P2 + 10¨HCl) =(2*-296.6 + 5*0) - (0 -95.4*10) = 360.8 kJ/mol

then... since this is possitive, this must NOT be spontanous

for...

dG° = -RT*ln(K)

K = exp(dG°/(-RT))

K = exp(dG°/(-RT)) = exp(-360800/(8.314*298)) = 5.6916*10^-64

which sttes it doest not favours producst, i.e. K < 1